Rapid River, Inc., has a 7.5% coupon bond that matures in 9 years. The bond pays interest semi-annually. What is the market price of a $1,000 face value bond if the yield to maturity is 6.8%?
Answer #1

Answer:

Given

face value of Bond F=$1000

Coupon rate =7.5%

Semiannual coupon payment C=7.5%*1000/2=$37.5

Years to maturity=9 years

Periods n=9*2=18 semi annual

Yield to maturity=6.8%

so r=6.7%/2=3.4%

So Price of Bond=C*(1-(1+r)^-n)/r + F/(1+r)^n

Price of bond=37.5*(1-(1+3.4%)^-18)/3.4% + 1000/(1+3.4%)^18=$1046.55